a rectangle pool has an area of 800 ft^2. the length is 10 feet longer than the width. find the dimensions of the pool. solve by completing the square. round answers to the nearest tenth of a foot.
Answer by Boreal(15235)

width=x
length=x+10
x(x+10)=800 ft^2
x^2 + 10 x= 800 ft^2.
x^2 +10x +25= 825
(x+5)^2=825
X+5+= +sqrt (825)
x=-5 +sqrt (825)=-5 +28.72=23.7 ft.
Check: 23.7ft width * 33.7ft length=798.69, which is in rounding error of 800 ft^2.